OpenStudy (anonymous):

find area of region y=xsin(x^2) on the interval [o,pi]

OpenStudy (anonymous):

u know u should integrate for finding area right?

OpenStudy (anonymous):

yeah but its been a minute since ive done this and im extremely rusty

OpenStudy (anonymous):

i know you use f(a) - f(b)

OpenStudy (anonymous):

emm..this is what we have\[\int_{0}^{\pi} x\sin x^2 \ \text{d}x\]

OpenStudy (anonymous):

can you show me the steps? is it simply f(a)-f(b) of x - f(a)-f(b) of sinx^2?

OpenStudy (anonymous):

I think u-sub works here, right?

OpenStudy (agent47):

He didn't say under the curve, so I'm assuming you'd need to break that integral up into 3 parts, and take the abs value of one part because it goes below x axis

OpenStudy (anonymous):

well this is a u-sub integration let\[u=x^2\]\[du=2x\text{d}x\]\[\int_{0}^{\pi^2} \frac{1}{2} \sin u \text{d}u\]

OpenStudy (anonymous):

ahh...thats a good point @Agent47 mentioned

OpenStudy (agent47):

@Agent47 always makes good points :D!

OpenStudy (anonymous):

its just find the area enclosed by that equation. honestly i dont know where to begin to solve

OpenStudy (anonymous):

but id say its just integrating because its impossible to solve \[x\sin x^2=0\]

OpenStudy (anonymous):

exactly. I was going to use integration by parts and set uprime to \[\sin^{2}\] and v = x , vprime=1

OpenStudy (anonymous):

by parts should work too,

OpenStudy (anonymous):

It's more difficult, though . .

OpenStudy (anonymous):

u-substitution is short and sweet.

OpenStudy (anonymous):

im not familiar with u-substitution

OpenStudy (anonymous):

Check @mukushla 's third reply; he sets it up for you.

OpenStudy (anonymous):

ohhhhhhhh

OpenStudy (anonymous):

Isn't that a lot prettier than an integration-by-parts mess?

OpenStudy (anonymous):

so completely bypass integration by parts and just go straight into solving the definite integral, subbing u out for pi and 0 and computing f(a) - f(b)???

OpenStudy (anonymous):

yeah much much prettier

OpenStudy (anonymous):

i still havent found the answer

OpenStudy (anonymous):

its not \[\frac{ 1 }{ 2 }-\cos \int\limits_{0}^{\pi}\] is it?

OpenStudy (anonymous):

im sorry cosu

OpenStudy (anonymous):

Ok, start with the substitution like Mukushla showed: Let u = x^2, so du = 2x dx

OpenStudy (anonymous):

x dx = half of du, so your integral becomes \[\int\limits x \sin(x^2) dx \rightarrow \frac{1}{2} \int\limits \sin(u) du\]

OpenStudy (anonymous):

the integral of sin(u) is -cos(u), so you have \[-\frac{1}{2} \cos (u), u=x^2 \rightarrow -\frac{1}{2} \cos (x^2)\] Evaluate from 0 to π and you're done.

OpenStudy (anonymous):

Be careful with the limits. I usually like to leave it indefinite until the end and then undo the substitution so I can keep the same limits, but really \[\int\limits_{0}^{π} x \sin(x^2) dx = \frac{1}{2}\int\limits_{0}^{π^2}\sin(u) du, \space u=x^2\]

OpenStudy (anonymous):

So the final answer should read... \[-\frac{ 1 }{ 2 }\cos(\pi ^{2}) -(-\frac{ 1 }{ 2 }\cos(0) )\]

OpenStudy (anonymous):

which gives -1/2 * 1 - (-1/2 * 1)???

OpenStudy (anonymous):

which is 0.. which cant be right

OpenStudy (anonymous):

Depends . . If you are subtracting areas below the x-axis, then maybe. If you add all the absolute values together, it should be something closer to 1.

OpenStudy (anonymous):

But the way I solved it is right though?

OpenStudy (anonymous):

Hmm, what's cos(π^2)?

OpenStudy (anonymous):

cos(pi) is -1, which squared would equal 1. My calculator read -.9026853619

OpenStudy (anonymous):

Square the π before taking cosine.

OpenStudy (anonymous):

(order of operations . . .)

OpenStudy (anonymous):

i still get the same answer

OpenStudy (anonymous):

You mean -.9026853619, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, then go from there.

OpenStudy (anonymous):

.951342681?

OpenStudy (anonymous):

Yeah, that's what my calculator is telling me too.

OpenStudy (anonymous):

Lol alright cool... thankyou

OpenStudy (anonymous):

You're welcome. I'm still curious if it was meant to subtract the areas below the x-axis or to add the absolute values of the areas.. I'm also confused why @mukushla said x*sin(x^2)=0 was impossible to solve; I found x=0, x=sqrt(π), x=sqrt(2π), and x=sqrt(3π) on that interval. Using those, you could set up four intervals of integration to find the positive areas of all the regions, . . but perhaps that's more work than it's worth.

OpenStudy (anonymous):

ahh...man my brain not functioning well these days .. u r right @CliffSedge .. its possible